Define a function A: N x N -> N as follows: A(m, n) ={2n, if m= 0; 0, if m≥1 and n= 0; 2, if m≥1 and n= 1; A(m-1, A(m, n-1)), if m≥1 and n≥2

Show that A(m, 2) = 4 whenever m≥1.
Hint: Induct on m.
1
Expert's answer
2020-12-13T19:28:04-0500

We shall prove by inducting over "m".

When "m=1" .

"A(1,2)=A(0,A(1,1))=A(0,2)=4."

It is true for "m=0".


Suppose it is true for some "m=k\geq1" . We shall prove that it holds for "m=k+1" .


"A(k+1,2)=A(k,A(k+1,1))\\"

Since "k\geq1, k+1\geq2" . Hence, "A(k+1,1)=2"

So,

"A(k+1,2)=A(k,2)"

We know that it is true for "m=k" . So, "A(k,2)=4" .


Therefore, "A(k+1,2)=4."

This shows that it holds for "m=k+1" .


Hence, "A(m,2)=4" for "m \geq1" .